OpenStudy (anonymous):

How do you solve -sin^2 x=2cosx-2

OpenStudy (anonymous):

\[-2\sin^2x=2cosx-2\]

zepdrix (zepdrix):

Adding 2sin^2x to each side gives us,\[\Large 0=2\sin^2x+2\cos x-2\]Divide both sides by 2,\[\Large 0=\sin^2x+\cos x-1\] From here we want to remember our `Square Identity`:\[\large \sin^2x+\cos^2x=1\qquad\to\qquad \color{royalblue}{\sin^2x=1-\cos^2x}\]Applying this identity to our problem,\[\Large 0=\color{royalblue}{\sin^2x}+\cos x-1\]Gives us,\[\Large 0=\color{royalblue}{1-\cos^2x}+\cos x-1\]

zepdrix (zepdrix):

Just requires a few more steps from here. Confused by any of that?

OpenStudy (anonymous):

no, but what do i do after getting\[0=cosx-\cos^2x\]

OpenStudy (anonymous):

o, is x=pi/2

zepdrix (zepdrix):

Factor cosx out of each term.\[\Large 0=\cos x(1-\cos x)\]Then by the Zero Factor Property we set each term equal to zero and solve separately.\[\Large 0=\cos x \qquad\qquad\qquad 0=1-\cos x\] x=pi/2 sounds correct, that is ONE of our solutions :)